OpenStudy (anonymous):

Identify the vertical asymptotes of f(x) = 10 over quantity x squared minus 7x minus 30

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

10 / (x^2-7x-30) wont the asymptote be formed by finding out what minimum value x can be and still be true in the equation: x^2 - 7x - 30, because we can't take the square of a negative number, right?

OpenStudy (anonymous):

\[10 \over x^2 - 7x - 30\]

OpenStudy (anonymous):

Yeah we can't square a negative

OpenStudy (anonymous):

also, these are my answer choices if it helps... x = 10 and x = 3 x = 10 and x = -3 x = -10 and x = 3 x = -10 and x = -3

OpenStudy (anonymous):

so take the equation part x^2 -7x -30 set it equal to zero, and solve for the roots make sense?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so what two number can we add to make -7 or p + q = -7 and what two numbers timesed by each other makes -30 or p*q = -30

OpenStudy (anonymous):

15(-2) and... -2(3.5)

OpenStudy (anonymous):

maybe if I just did it this way it'd be easier for you (x-10)(x+ ? ) = x^2 -7x -30

OpenStudy (anonymous):

x + 10?

OpenStudy (anonymous):

if we foiled: (x-10)(x+10) we'd get: x^2 -100 that dones't equal x^2 - 7x -30 that we need. try one more time, then i'll step in ^_^

OpenStudy (anonymous):

hmm... 3 lol?

OpenStudy (anonymous):

haha ya! (x-10)(x+3) so then we find the roots by this method thing... here: (x-10) = 0, solve for x (x+3) = 0, solve for x

OpenStudy (anonymous):

x = 10 and x = -3

OpenStudy (anonymous):

yep! those are the vertical asytompes

OpenStudy (anonymous):

Thanks!